OpenStudy (anonymous):

if you had a problem that looked like this: 4,500=4,200+3%c how would you get the c by it's self

OpenStudy (anonymous):

do I subtract

OpenStudy (texaschic101):

4500 = 4200 + .03c (3% = .03) 4500 - 4200 = .03c 300 = .03c (divide by .03 to get c by itself) Can you take it from here ?

OpenStudy (anonymous):

4,500-3=4,200+3%c-3 would it look like this

OpenStudy (texaschic101):

you do not subtract the .03 .... you divide by it.

OpenStudy (anonymous):

oh okay I am confussed about how it should look after I do that

OpenStudy (texaschic101):

Let me show you the whole problem.... 4500 = 4200 + 3%c (turn the 3% into .03) 4500 - 4200 = 4200 + .03c - 4200 (subtract 4200 from both sides) 300 = .03c (now divide by .03 to get c by itself) 300/.03 = .03c/.03 10,000 = c check... 4500 = 4200 + .03c (sub 10,000 in for c) 4500 = 4200 + .03(10,000) 4500 = 4200 + 300 4500 = 4500 (correct) Do you understand ?

OpenStudy (anonymous):

yeah thank you

OpenStudy (texaschic101):

you are very welcome...thats what I am here for :)